40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Mon Feb 23 14:57:23 2015, Zach, Laser, M2 ISS, 3-mm diode initial noise measurement 3mm_diode_noise_5V.png
    Reply  Mon Feb 23 17:57:49 2015, rana, Laser, M2 ISS, 3-mm diode initial noise measurement 
       Reply  Mon Feb 23 19:55:08 2015, Koji, Laser, M2 ISS, 3-mm diode initial noise measurement 
       Reply  Thu Feb 26 18:47:59 2015, Zach, Laser, M2 ISS, 3-mm diode dark current and noise 7845_turnon.PNG7843_turnon.PNG3mm_pd_noise_5V_2_25_25.png
Message ID: 1212     Entry time: Thu Feb 26 18:47:59 2015     In reply to: 1208
Author: Zach 
Type: Laser 
Category: M2 ISS 
Subject: 3-mm diode dark current and noise 

Given that we see some excess noise in our 3-mm Laser Components diodes (IG17X3000G1i), especially with one of them, I went ahead and did a more careful measurement of both the dark currents and noise.

To make this measurement, I switched to OPA140 transimpedance amps on the M2 readout board and used a 1-M\Omega transimpedance.

Dark current

The OPA140 has a bias/offset current of 10 pA max and an offset voltage of 120 uV max, the latter which therefore limits the DC current sensitivity with this transimpedance to 0.12 nA. There is also some allowed current variation over temperature (±3 nA bias and ±1 nA offset over -40 to +125 °C), so this can add some more DC uncertainty if the lab temperature is a few degrees away from 25 °C. Plugging the outputs into the DMM without the diodes connected, I measured 0.0 mV and 0.2 mV on amps 1 and 2, respectively. This is consistent with the op amp spec.

The Laser Components spec for the dark current (at 5-V bias, where I measured it) is 20 nA typ, 100 nA max. Plugging in the diodes while keeping them in a blocked box and with the room lights off, I measured the following bias currents (output voltage divided by 1 M\Omega):

  • S/N 7842: 78.9 nA
  • S/N 7843: 61.7 nA
  • S/N 7844: 43.8 nA
  • S/N 7845: > 200 nA (started below 100 nA, increased continuously while energized---see below)

The first 3 diodes are within the max spec, while 7845 seems to be exhibiting some catastrophic failure mode where the dark current is avalanching whenever the bias is engaged. Below is a plot of the measured amplifier output after a turn-on of this diode, with one of a healthy diode for comparison to the right. This was taken in the middle of the testing, and the last measurement of the current before this turn-on was around 140 nA. As you can see, there is an initial slew (not inconsistent with the timescale of the bias turn-on), followed by a slow but monotonic increase of the dark current over time. When this was repeated, the initial slew brought the current again to the last-known highest level.

 

So, as you can see, S/N 7845 is clearly broken. Maybe we can get a replacement

Noise

I used the same transimpedance amp setup to measure the noise. All diodes show spectactularly higher noise than the advertised level of 3.2 x 10-14 W/rtHz NEP (~3 x 10-14 A/rtHz), with a 1/f characteristic that, if extrapolated, would not intercept the quoted spec until ~1 MHz. A frequency for this spec is not mentioned on the datasheet. In all cases, the circuit was allowed to equilibrate for a few minutes before a measurement was made. The spectra below were found to be stationary, with the exception of occasional glitches.

The readout noise is limited from several 100 mHz up to near 1 kHz by the Johnson noise of the 1-M\Omega transimpedance resistor, above which there is some noise peaking centered around 40 kHz that is not inconsistent with other measurements I have made with this op amp in very-high-impedance environments (c.f., 40m:8151).

What gives?

Quote:

IF the DC dark current is out of spec, we might be able to get a replacement. Might be specs on the website. I think Frank had a Keithley instrument to measure dark currents that are low - probably in his diode destruction elogs or DCC docs.

 

ELOG V3.1.3-